OpenStudy (dtan5457):

If z varies directly with he product of x and y, then z is said to vary __________ with x and y.

OpenStudy (dtan5457):

@Nnesha @DanJS

OpenStudy (dtan5457):

Is this direct variation?

OpenStudy (danjs):

\[z~~\alpha~~x*y\]

OpenStudy (danjs):

So z = (constant)*x*y z = k*x*y

OpenStudy (dtan5457):

so directly?

OpenStudy (jordanloveangel):

@dtan5457 go here http://www.shelovesmath.com/algebra/beginning-algebra/direct-inverse-and-joint-variation/ this helps me alot

OpenStudy (danjs):

I would assume, but the question already says ' z varies directly with he product of x and y' so is there some other special word they call it?

OpenStudy (danjs):

Joint variation is just like direct variation, but involves more than one other variable.

OpenStudy (dtan5457):

Not sure...just started learning this today

OpenStudy (danjs):

Joint

OpenStudy (danjs):

jointly i guess

OpenStudy (danjs):

z varies JOINTLY with x and y

OpenStudy (danjs):

found it on that site jordan linked, didn't know there was a special name for that

OpenStudy (dtan5457):

Can you quickly explain the difference between jointly and directly?

OpenStudy (danjs):

Joint is where the dependent variable (z) varies directly with more than 1 independent variable, here (x and y)

OpenStudy (danjs):

same as direct variation, just more than 1 independent variable

OpenStudy (dtan5457):

so z=x (direct) z=xy (jointly)

OpenStudy (dtan5457):

direct=1 variable

OpenStudy (danjs):

Direct Variation \[z~~\alpha~~x\] \[z = k*x\] Joint Variation is a form of direct variation, but z will be directly proportional to more than 1 variable, x and y.

OpenStudy (danjs):

\[z~~\alpha~~x*y\] \[z = k*(x*y)\]

OpenStudy (danjs):

sorry the site server is bogged down or something, lagging like crazy

OpenStudy (danjs):

Not z = x \[z~~\alpha~~~x\] z is proportional to x, you need a constant of proportionality to make an equation z = k*x

OpenStudy (danjs):

Z varies Directly with X --- z = k*x Z varies Jointly with x and y ----z = k*x*y Joint is just a fancy word meaning z varies directly with more than 1 other variable

Nnesha (nnesha):

combined variation 0r joined variation for example jz varies jointly with x and y ^^^^ \(\huge\color{green}{{\rm z }=k xy ~}\) but remember jointly is always in applied direct relation so if there is a word "jointly" its mean direct

Nnesha (nnesha):

uh-oh lost the connection....trying to reconnect now. again again agian